OpenStudy (anonymous):

Put the equation in the form ax^2+bx+c=0 3x^2+4x=(3x-1)^2-2x

OpenStudy (anonymous):

3x^2+4x=(3x-1)^2-2x 3x^2+4x=(9x^2+1 -6x) -2x 3x^2+4x- (9x^2+1 -6x) + 2x =0 -6x^2 +12x -1 = 0